# Wukong: Towards a Scaling Law for Large-Scale Recommendation

### 04 Mar 2024 v1

@@ Abstract
We conducted extensive evaluations on six public datasets, and our results demonstrate that Wukong consistently outperforms state-of-the-art models quality-wise. Further, we assessed Wukong’s scalability on an internal, large-scale dataset. The results show that Wukong retains its superiority in quality over state-of-the-art models, while holding the scaling law across two orders of magnitude in model complexity, extending beyond 100 GFLOP/example or equivalently up to GPT-3/LLaMa-2 scale of total training compute, where prior arts fall short.
@@
Figure 1: Wukong outperforms existing state-of-the-art models while demonstrating a scaling law in the recommendation domain across two orders of magnitude in model complexity, extending beyond 100 GFLOP/example or equivalently up to GPT-3/LLaMa-2 scale of total training compute, where prior arts fall short.
@@ 1 Introduction
We evaluated Wukong’s performance using six public datasets and a large-scale internal dataset. The results demonstrate that Wukong outperforms state-of-the-art models across all public datasets in terms of AUC, indicating the effectiveness of Wukong’s architecture and its ability to generalize across a wide range of recommendation tasks and datasets. On our internal dataset, Wukong not only significantly outperforms existing models in terms of quality at comparable levels of complexity but also shows continuous enhancements in quality when scaled up across two orders of magnitude in model complexity, extending beyond 100 GFLOP/example or equivalently up to GPT-3/LLaMa-2 scale of total training compute, where prior arts fall short.
@@ 7.2 Results
Quality vs. Compute Complexity    In Fig. 1, we depict the relationship between quality and compute complexity. The results show that Wukong consistently outperforms all baselines across various complexity levels, achieving over 0.2% improvement in LogLoss. Notably, Wukong holds its scaling law across two orders of magnitude in model complexity – we observe a significant 0.4% enhancement in quality with a 200-fold complexity rise , up to GPT-3/LLaMa-2 scale in total training compute 111 Projecting the 160 PF-days of total training compute on the internal dataset to a 365-day timeframe, assuming the use of continuous online training. approximately translating to a 0.1% improvement for every quadrupling of complexity. Among baselines, AFN+, DLRM and FinalMLP tend to reach a plateau after a certain complexity level, while AutoInt+, DCNv2 and MaskNet failed to further enhance quality 222 AutoInt+ and DCNv2 consistently faced significant training instability issue when further scaled up. AutoInt+ recovered from loss explosion, albeit with reduced model quality; while DCNv2 failed to recover, and its quality was estimated from performance before the explosion. MaskNet was hindered by excessive memory consumption, leading to out-of-memory errors, blocking further scaling up.. Nonetheless, even DCNv2, the top-performing baseline, demands a 40-fold increase in complexity to match Wukong’s quality.

### 08 Mar 2024 v2

@@ Abstract
We conducted extensive evaluations on six public datasets, and our results demonstrate that Wukong consistently outperforms state-of-the-art models quality-wise. Further, we assessed Wukong’s scalability on an internal, large-scale dataset. The results show that Wukong retains its superiority in quality over state-of-the-art models, while holding the scaling law across two orders of magnitude in model complexity, extending beyond 100 GFLOP/example or equivalently up to Large Launguage Model (GPT-3) training compute scale, where prior arts fall short.
@@
Figure 1: Wukong outperforms existing state-of-the-art models while demonstrating a scaling law in the recommendation domain across two orders of magnitude in model complexity, extending beyond 100 GFLOP/example or equivalently up to Large Launguage Model (GPT-3) training compute scale, where prior arts fall short.
@@ 1 Introduction
We evaluated Wukong’s performance using six public datasets and a large-scale internal dataset. The results demonstrate that Wukong outperforms state-of-the-art models across all public datasets in terms of AUC, indicating the effectiveness of Wukong’s architecture and its ability to generalize across a wide range of recommendation tasks and datasets. On our internal dataset, Wukong not only significantly outperforms existing models in terms of quality at comparable levels of complexity but also shows continuous enhancements in quality when scaled up across two orders of magnitude in model complexity, extending beyond 100 GFLOP/example or equivalently up to 2336 PF-days of total training compute, normalized to 365 days assuming online training (Zhai et al., 2024), comparable to Large Launguage Model (GPT-3, 3640PF-days) training compute scale (Brown et al., 2020), where prior arts fall short.
@@ 7.2 Results
Quality vs. Compute Complexity    In Fig. 1, we depict the relationship between quality and compute complexity. The results show that Wukong consistently outperforms all baselines across various complexity levels, achieving over 0.2% improvement in LogLoss. Notably, Wukong holds its scaling law across two orders of magnitude in model complexity – approximately translating to a 0.1% improvement for every quadrupling of complexity. Among baselines, AFN+, DLRM and FinalMLP tend to reach a plateau after a certain complexity level, while AutoInt+, DCNv2 and MaskNet failed to further enhance quality 111 AutoInt+ and DCNv2 consistently faced significant training instability issue when further scaled up. AutoInt+ recovered from loss explosion, albeit with reduced model quality; while DCNv2 failed to recover, and its quality was estimated from performance before the explosion. MaskNet was hindered by excessive memory consumption, leading to out-of-memory errors, blocking further scaling up.. Nonetheless, even DCNv2, the top-performing baseline, demands a 40-fold increase in complexity to match Wukong’s quality.